In an economy with no taxes and no imports, disposable income decreases from $6,000 to $4,000. If consumption decreases from $4,500 to $3,000, the marginal propensity to save is:


can someone show me the math to this question
Answer #1

MPS= 1- MPC

MPS= marginal propensity to save

MPC= marginal propensity to consume

MPC= change in consumption / change in income OR ▲C/ ▲Y

= 3000- 4500 / 4000-6000

= - 1500/ - 2000

= 0.75

MPS= 1- 0.75

= 0.25